Solve the following based off the following data:

6) Given 1 Find P37 2) Five point summary 3) Box and whisker plot 4 Is this data normal? 6 Hrowwnn e Wap funt vt Noore

1) Find P37

2) Five Point Summary

3) Box and Whisker Plot

4) Is this data normal?
Answer #1

Solution(1)

P37=37/100*(27+1)

= 10.36

P37=10 th value+0.36*(11th value-10th value)

=55+0.36*(56-55)

= 55.36

P37=55.36

Solution-(2)

MIN= 35.0

Q1=51.0

Q2=64.0

Q3=84.5

MAX=99.0

Solution-3:
Rcode:

score <- c(35,37,39,42,46,48,49,
53,55,55,56,57,
63,64,68,
70,71,72,73,
84,85,87,
91,93,95,97,99)
boxplot(score,horizontal = TRUE,col="green")

40 5060708090 100

In boxplot center line is median=64=Q2

min=35

Q1=51

Q3=84.5

max=99

Soluiton:

mean=sum of values/total values

=1784/27

= 66.07407

median=middlemost value after sorting a number

35 37 39 42 46 48 49 53 55 55 56 57 63 64 68 70 71 72 73 84 85 87 91 93 95 97 99

median=14th item=64

mean>median

Positively skewed

NOT NORMAL